# Stable initial conditions and analytically approximate solutions of cosmological perturbations in a modified loop quantum cosmology

### v2

@@ IV Mode Functions in Different Phases
where $\lambda\;(\equiv k^{-1})$ denotes the comoving wavelength of the mode $k$. Note that such a defined quantity $\lambda_{H}^{2}$ becomes negative when the effective mass is positive. In Fig. 6, we plot $\lambda_{H}^{2}$ schematically for the Starobinsky potential, where the moments $t_{H}$ and $t_{i}$ are defined, respectively, by $a^{\prime\prime}(t_{H})=a^{\prime\prime}(t_{i})=0$, so $t_{i}$ represents the beginning of the inflationary phase, and during the slow-roll inflation (Region III), we have $\lambda_{H}^{2}\approx L_{\text{H}}^{2}/2\simeq 1/(2a^{2}H^{2})$, which is exponentially decreasing, and all the modes observed today were inside the comoving Hubble radius at $t=t_{i}$. Between the times $t_{H}$ and $t_{i}$, $\lambda_{H}^{2}$ is negative, and $\Omega^{2}_{\text{tot}}$ is strictly positive. Therefore, during this period the mode functions are oscillating, while during the epoch between $t_{B}$ and $t_{H}$, some modes ($k^{-2}>k_{B}^{-2}$) are inside the comoving Hubble radius, and others ($k^{-2}<k_{B}^{-2}$) are outside it right after the bounce, where $k_{B}\equiv\lambda_{B}^{-1}(t_{B})$. In the contracting phase, when $t\ll t_{B}$, the universe is quasi-de Sitter and $\lambda_{H}^{2}\simeq 1/(2a^{2}H^{2})$ increases exponentially toward the bounce $t\rightarrow t_{B}$, as now $a(t)$ is decreasing exponentially. However, several Planck seconds before the bounce, the universe enters a non-de Sitter state, during which $\lambda_{H}^{2}$ starts to decrease until the bounce, at which a characteristic Planck scale $k_{B}\;(\equiv 1/\lambda_{H})$ can be well defined. Therefore, for $t\ll t_{B}$, some modes are outside the comoving Hubble radius, and become non-adibatic. Then, it is impossible to impose the BD vacuum. To study the perturbation further, let us consider the modified MS equation in each of these phases separately.
@@ V Analytical Solutions of the Mode Function
where $\dot{y}\equiv dy/d\xi$, and ${\cal{F}}(\xi)$ is a function of $\xi$ only, the choice of which depends on the properties of the turning points (or roots) of $g(y)=0$ [67, 46]. In terms of $U(\xi)$ Eq.(62) reads
@@ V.2 $k_{e}\lesssim k\lesssim k_{2}$
$\mathfrak{D}(b,x)\equiv\frac{1}{2}x\sqrt{\frac{1}{4}x^{2}-b}-b\ln\left(\frac{\frac{1}{2}x+\sqrt{\frac{1}{4}x^{2}-b}}{\sqrt{b}}\right)+\frac{\pi}{4}+\phi(b),$ (101)
$\phi(b)\equiv\frac{b}{2}-\frac{b}{4}\ln{b^{2}}+\frac{1}{2}\text{ph}\Gamma\left(\frac{1}{2}+ib\right),$
@@ VII Conclusions
In this paper, we applied the UAA method [46] to find the first-order approximation solution of the mode function that satisfies the generalized MS equation (34) in the dressed metric approach within the framework of mLQC-I [29, 28]. Thanks to the universal and analytical background solutions given by Eqs.(II.2), (28) and (31), which led to such analytical analyses possible.
In particular, with the extension of the effective potential across the quantum bounce into the deep remote contracting phase considered in [42], we analyzed the effective mass function $m_{\text{eff}}$ defined by Eqs.(42) and (43) in detail. As shown in Fig. 6, the effective mass squire is always negative, so the effective Hubble horizon $\lambda_{H}^{2}\equiv-m_{\text{eff}}^{-2}$ is positive and decreases exponentially as $t\rightarrow$-$\infty$ [cf. Eqs.(48) and (49)], while in the post-bounce phases, $\lambda_{H}$ exhibits complicated behaviors. Dividing the whole evolution into four different phases, pre-de Sitter, (extended) bouncing, transition and inflation, the mode function is given analytically by Eq.(53) in the pre-de Sitter phase, and Eq.(58) in the transition and inflationary phases. Therefore, the mode function is not known only during the bouncing phase.
@@
In the forthcoming work, we shall calculate $f(k)$ explicitly, after generalizing the current firs-order approximation to high-order ones, so that the approximate solutions will have enough accuracy [68, 70], in order to compare our theoretical predictions with current and forthcoming observations [52]. To have the mLQC-I model physically viable, several other important issues also need to be addressed, including particle creation, anisotropy and non-Gausianity [18]. We wish to come back to these issues soon.
@@ Acknowledgments
We thank Prof. Robert Brandenberger for reading our manuscript carefully, and valuable discussions, suggestions and comments. We also thank Dr. Bao-Fei Li for the early stage collaboration of the project and Drs. Qiang Wu and Tao Zhu for their kind help, suggestions and comments. R.P. and A.W. are partially supported by the US NSF grant, PHY-2308845, while J.S. is supported by Baylor through the Baylor Physics graduate program.

### v3

@@ IV Mode Functions in Different Phases
where $\lambda\;(\equiv k^{-1})$ denotes the comoving wavelength of the mode $k$. Note that such a defined quantity $\lambda_{H}^{2}$ becomes negative when the effective mass is positive. In Fig. 6, we plot $\lambda_{H}^{2}$ schematically for the Starobinsky potential, where the moments $t_{H}$ and $t_{i}$ are defined, respectively, by $\ddot{a}(t_{H})=\ddot{a}(t_{i})=0$, so $t_{i}$ represents the beginning of the inflationary phase, and during the slow-roll inflation (Region III), we have $\lambda_{H}^{2}\approx L_{\text{H}}^{2}/2\simeq 1/(2a^{2}H^{2})$, which is exponentially decreasing, and all the modes observed today were inside the comoving Hubble radius at $t=t_{i}$. Between the times $t_{H}$ and $t_{i}$, $\lambda_{H}^{2}$ is negative, and $\Omega^{2}_{\text{tot}}$ is strictly positive. Therefore, during this period the mode functions are oscillating, while during the epoch between $t_{B}$ and $t_{H}$, some modes ($k^{-2}>k_{B}^{-2}$) are inside the comoving Hubble radius, and others ($k^{-2}<k_{B}^{-2}$) are outside it right after the bounce, where $k_{B}\equiv\lambda_{B}^{-1}(t_{B})$. In the contracting phase, when $t\ll t_{B}$, the universe is quasi-de Sitter and $\lambda_{H}^{2}\simeq 1/(2a^{2}H^{2})$ increases exponentially toward the bounce $t\rightarrow t_{B}$, as now $a(t)$ is decreasing exponentially. However, several Planck seconds before the bounce, the universe enters a non-de Sitter state, during which $\lambda_{H}^{2}$ starts to decrease until the bounce, at which a characteristic Planck scale $k_{B}\;(\equiv 1/\lambda_{H})$ can be well defined. Therefore, for $t\ll t_{B}$, some modes are outside the comoving Hubble radius, and become non-adibatic. Then, it is impossible to impose the BD vacuum. To study the perturbation further, let us consider the modified MS equation in each of these phases separately.
@@ V Analytical Solutions of the Mode Function
where $\dot{y}\equiv dy/d\xi$  [^5] It should be noted that in this paper we also use an over dot to denote the derivative with respect to the cosmic time $t$, for example, $\dot{a}\equiv da/dt,\dot{\rho}\equiv d\rho/dt,\dot{\phi}=d\phi/dt$, etc. However, the dot over $y$ in this paper always represents the derivative of $y$ with respect to $\xi$, $\dot{y}\equiv dy/d\xi$. This is the only exception, and in all other cases an over dot always denote the derivative with respect to the cosmic time $t$, without causing any confusion., and ${\cal{F}}(\xi)$ is a function of $\xi$ only, the choice of which depends on the properties of the turning points (or roots) of $g(y)=0$ [94, 76]. In terms of $U(\xi)$ Eq.(62) reads
@@ V.2 $k_{e}\lesssim k\lesssim k_{2}$
$\mathfrak{D}(b,x)\equiv\frac{1}{2}x\sqrt{\frac{1}{4}x^{2}-b}-b\ln\left(\frac{\frac{1}{2}x+\sqrt{\frac{1}{4}x^{2}-b}}{\sqrt{b}}\right)+\frac{\pi}{4}+\phi(b),$
$\phi(b)\equiv\frac{b}{2}-\frac{b}{4}\ln{b^{2}}+\frac{1}{2}\text{ph}\Gamma\left(\frac{1}{2}+ib\right),$ (101)
@@ VII Conclusions
In this paper, we applied the UAA method [76] to find the first-order approximation solution of the mode function that satisfies the generalized MS equation (34) in the dressed metric approach within the framework of mLQC-I [89, 62]. Thanks to the universal and analytical background solutions given by Eqs.(II.2), (28) and (31), which led to such analytical analyses possible  [^8] It should be noted that the universal behavior of the background evolution holds only for initially kinetic-energy-dominated conditions at the bounce, $\dot{\phi}_{B}^{2}\gg 2V(\phi_{B})$. For other initial conditions, such as initially potential-energy-dominated ones, such universality gets lost and inflationary phase may not occur [58, 59, 60]. This is true not only in mLQC-I but also in LQC and other models [62, 33, 96]. However, since inflation happens generically not only in LQC [21] but also in mLQC-I and other models [60], without loss of the generality, in this paper we have focused on the initial conditions that lead generically to inflation in the post-bounce region..
In particular, with the extension of the effective potential across the quantum bounce into the deep remote contracting phase considered in [57], we analyzed the effective mass function $m_{\text{eff}}$ defined by Eqs.(42) and (43) in detail. As shown in Fig. 5, the effective mass square is always negative, so the effective Hubble horizon $\lambda_{H}^{2}\equiv-m_{\text{eff}}^{-2}$ is positive and decreases exponentially as $t\rightarrow$-$\infty$ [cf. Eqs.(48) and (49)], while in the post-bounce phases, $\lambda_{H}$ exhibits complicated behaviors. Dividing the whole evolution into four different phases, pre-de Sitter, (extended) bouncing, transition and inflation, the mode function is given analytically by Eq.(53) in the pre-de Sitter phase, and Eq.(58) in the transition and inflationary phases. Therefore, the mode function is not known only during the bouncing phase.
@@
The function $f(k)$ certainly depends on the choice of the initial conditions. In Sec. VI, using a method developed by Birrell and Davies [28], we identify an initial state in the remote contracting phase given by Eq.(162), which turns out to be stable, minimize particle creations [78], and diagonalize the Hamiltonian [47], despite the fact that at this time some modes may be still outside of the Hubble horizon and not in their adiabatic states. It would be also very interesting to compare such obtained $f(k)$ with the ones obtained by other initial conditions [4, 55, 54, 17, 18, 43, 39, 68, 74, 71, 13].
In the forthcoming work, we shall calculate $f(k)$ explicitly, after generalizing the current firs-order approximation to high-order ones, so that the approximate solutions will have enough accuracy [93, 98], in order to compare our theoretical predictions with current and forthcoming observations [1].
To have the mLQC-I model physically viable, several other important issues also need to be addressed, including particle creation, anisotropy and non-Gausianity [9]. We wish to come back to these issues soon.
Finally, we note that in the studies of cosmological perturbations in LQC/mLQCs, there exist two main approaches in the literature, the dressed metric [2, 3, 4], and the hybrid [48]. In the current paper, we have mainly focused ourselves on the dressed metric approach. It would be very interesting to extend such studies to the hybrid approach, of which we just studied numerically in [57] and observed some differences. It is important to understand such differences analytically, and then compare them with observations.
@@ Acknowledgments
We thank Prof. Robert Brandenberger for reading our manuscript carefully, and valuable discussions, suggestions and comments. We also thank Dr. Bao-Fei Li for the early stage collaboration of the project and Drs. We also thank Qiang Wu and Tao Zhu for their kind help, suggestions, and comments. R.P. and A.W. are partially supported by the US NSF grant, PHY-2308845, while J.S. is supported by Baylor through the Baylor Physics graduate program.